5. Use the sign of the derivative to establish the inequaltiy ln(1 + x) > x − x 2 2 , ∀ x > 0


1
Expert's answer
2021-04-29T16:15:52-0400

To establish this inequality we have to made a assumption first.

"f(x) = ln(1+x) -x+x^2"

Hence,

"f'(x ) = \dfrac{1}{1+x}-1+2x"


"f'(x) = \dfrac{x+2x^2}{1+x}"


For "x>0" , We get "f'(x) >0"


Hence, we clearly say that "f(x)>f(0)"


Hence, "ln(1+x)-x+x^2>0"


OR


"ln(1+x)>x-x^2" (Proved).